how many pounds of 40 cent nuts and how many pounds of 25 cent nuts must you a dealer use to produce a misture of 90 pounds to sell at 31 cents per pound?
Answer by stanbon(75887)

how many pounds of 40 cent nuts and how many pounds of 25 cent nuts must you a dealer use to produce a misture of 90 pounds to sell at 31 cents per pound?
---------
Equations:
Quantity: x + y = 90 lbs
Value:::40x + 25y = 31*90 cents
----------------------------------
Multiply Quantity by 40
----
40x + 40y = 40*90
---
Subtract from Value and solve for "y":
-15y = -9*90
y = 54 lbs (amt. of 25 cent nuts needed in the mixture)
---
Since x+y = 90, x = 36 lbs (amt. of 40 cent nuts needed in the mixture)
------
Cheers,
Stan H.